What is the wavelength, in nm, of radiation which has an energy of 216.0 kJ per mole of photons? mole of photons?
Answer #1

We know that

E = h v

Where,

E is energy,

ν is frequency

h is Plank's constant (h= 6.626 × 10-34 J s)

Using the value of energy provided in question.

E = 216.0 KJ = 216000 J

v = E/h = 216000 J / 6.626 × 10-34 J s = 32.598 x 1037 s-1

It is also known that ν = c / λ

where

ν is frequency

c = speed of light (3 x 108 m/sec)

So, λ = c / ν

by using the value of v we will get,

λ = (3 x 108 m/sec) / ( 32.598 x 1037 s-1)

λ = 9.203 X 10-31 m.